# Griffin: Towards a Graph-Centric Relational Database Foundation Model

### v1

@@ Appendix F Raw Results
Model Avg. Rank Seznam/charge Seznam/prepay Airbnb/destination Amazon/churn Diginetica/ctr Outbrain/ctr Rel-avito/user-clicks Rel-avito/user-visits
Sage 5.250 0.7917 0.8768 0.8424 0.7358 0.7273 0.6239 0.4916 0.5642
Gat 5.917 0.8053 0.8954 0.8428 0.7410 0.6741 0.6146 0.3694 0.6133
Pna 6.083 0.8000 0.8924 0.8464 0.7645 0.7011 0.6249 0.5550 0.6422
Hgt 5.583 0.7965 0.8805 0.8252 0.7551 0.6733 0.6260 0.5445 0.5451
DFS+MLP 6.333 0.7554 0.8248 0.7643 0.6815 0.6944 0.5456 0.5839 0.6521
DFS+Deepfm 7.250 0.7016 0.8092 0.8007 0.6667 0.7341 0.5289 0.5912 0.5844
DFS+Fttransformer 5.625 0.7473 0.8162 0.7863 0.6765 0.7412 0.5360 0.6247 0.6576
DFS+XGB 6.500 0.7600 0.8453 0.7561 0.6922 0.7219 0.5421 0.6028 0.6568
Griffin-unpretrained 3.500 0.7998 0.8941 0.8615 0.7307 0.7157 0.6246 0.6639 0.6261
Griffin-pretrained 2.833 0.8133 0.9058 0.8681 0.7417 0.7181 0.6253 0.6330 0.6468
@@
Sage 0.7194 0.8048 0.6767 0.6111 0.8470 0.8558 0.8861 0.6610 -0.9639
Gat 0.7048 0.7987 0.5011 0.5941 0.8284 0.8645 0.8853 0.6498 -0.9563
Pna 0.6874 0.7699 0.5007 0.5680 0.8366 0.8664 0.8896 0.6684 -0.9615
Hgt 0.7108 0.7064 0.5000 0.6457 0.8495 0.8670 0.8817 0.7119 -0.9636
DFS+MLP 0.7134 0.7796 0.6802 0.6518 0.8181 0.8326 0.8783 0.6772 -0.9847
DFS+Deepfm 0.6705 0.8133 0.6801 0.6210 0.8182 0.8212 0.8821 0.6315 -0.9946
DFS+Fttransformer 0.7302 0.8250 0.6800 0.6578 0.8034 0.8376 0.8749 0.6660 -0.9888
DFS+XGB 0.7158 0.8022 0.6786 0.6521 0.7906 0.8251 0.8675 0.7142 -0.9972
Griffin-unpretrained 0.7052 0.7855 0.6847 0.6722 0.9512 0.8457 0.8956 0.6680 -0.6117
Griffin-pretrained 0.7091 0.7795 0.6804 0.6908 0.9643 0.8435 0.8962 0.6978 -0.5994
@@
Sage -0.7239 -0.6276 -1.0984 -0.9205 -1.4130 -0.6151 -2.3820
Gat -0.9096 -0.6164 -1.2022 -0.9296 -1.4455 -0.6191 -2.3880
Pna -1.1861 -0.6387 -2.9402 -0.9649 -1.6745 -0.7015 -2.3840
Hgt -0.6888 -0.5588 -1.7845 -0.9777 -1.6348 -0.6126 -2.3830
DFS+MLP -0.6610 -0.6146 -1.2518 -0.9462 -2.1104 -0.6185 -2.3840
DFS+Deepfm -1.1784 -2.5796 -1.3540 -0.9753 -2.0998 -0.6118 -2.3810
DFS+Fttransformer -0.6339 -0.5934 -0.9256 -0.9793 -1.6947 -0.6460 -2.3830
DFS+XGB -0.6743 -0.6130 -1.2957 -0.9514 -2.1859 -0.6663 -2.3890
Griffin-unpretrained -0.6593 -0.5586 -0.8879 -0.7926 -1.1700 -0.5684 -2.3930
Griffin-pretrained -0.6586 -0.5694 -0.8962 -0.7945 -1.2148 -0.5518 -2.3890
@@
Task diginetica-downsample-ctr rel-hm-item-sales rel-hm-user-churn retailrocket-cvr seznam-charge seznam-prepay
@@
Task amazon-churn amazon-rating outbrain-small-ctr rel-avito-ad-ctr rel-avito-user-clicks rel-avito-user-visits
@@
Task rel-f1-driver-dnf rel-f1-driver-position rel-f1-driver-top3 stackexchange-churn stackexchange-upvote virus-wnv-pred
@@
Task airbnb-destination rel-trial-site-success rel-trial-study-adverse rel-trial-study-outcome talkingdata-demo-pred telstra-severity
@@
Task diginetica-downsample-ctr rel-hm-item-sales rel-hm-user-churn retailrocket-cvr seznam-charge seznam-prepay
@@
Task amazon-churn amazon-rating outbrain-small-ctr rel-avito-ad-ctr rel-avito-user-clicks rel-avito-user-visits
@@
Task rel-f1-driver-dnf rel-f1-driver-position rel-f1-driver-top3 stackexchange-churn stackexchange-upvote virus-wnv-pred
@@
Task airbnb-destination rel-trial-site-success rel-trial-study-adverse rel-trial-study-outcome talkingdata-demo-pred telstra-severity
@@
Task diginetica-downsample-ctr rel-hm-item-sales rel-hm-user-churn retailrocket-cvr seznam-charge seznam-prepay
@@
Task amazon-churn amazon-rating outbrain-small-ctr rel-avito-ad-ctr rel-avito-user-clicks rel-avito-user-visits
@@
Task rel-f1-driver-dnf rel-f1-driver-position rel-f1-driver-top3 stackexchange-churn stackexchange-upvote virus-wnv-pred
@@
Task airbnb-destination rel-trial-site-success rel-trial-study-adverse rel-trial-study-outcome talkingdata-demo-pred telstra-severity
@@
Task diginetica-downsample-ctr rel-hm-item-sales rel-hm-user-churn retailrocket-cvr seznam-charge seznam-prepay
@@
Task amazon-churn amazon-rating outbrain-small-ctr rel-avito-ad-ctr rel-avito-user-clicks rel-avito-user-visits
@@
Task rel-f1-driver-dnf rel-f1-driver-position rel-f1-driver-top3 stackexchange-churn stackexchange-upvote virus-wnv-pred
@@
Task airbnb-destination rel-trial-site-success rel-trial-study-adverse rel-trial-study-outcome talkingdata-demo-pred telstra-severity

### v2

@@ B.2 Metric Adjustments for Regression Tasks
B.3 Baseline Implementation Details
The baseline models are primarily adapted from the 4DBInfer framework, as most lacked a native implementation in RelBench. For the 4DBInfer datasets, we report the results directly from the original publication. For the RelBench datasets, while the Sage results are taken from the original report, all other baselines were re-evaluated. This process involved processing the RelBench datasets through the 4DBInfer pipeline. To ensure methodological consistency, we adhered to RelBench’s hyperparameter tuning strategy and incorporated a ResNet architecture into the encoder as specified in their design.
@@ Appendix F Raw Results
Dataset/Task Overall Score Seznam/charge Seznam/prepay Airbnb/destination Amazon/churn Diginetica/ctr Outbrain/ctr Rel-avito/user-clicks Rel-avito/user-visits
Metric Rank $\downarrow$ Accuracy $\uparrow$ Accuracy $\uparrow$ ROCAUC $\uparrow$ ROCAUC $\uparrow$ ROCAUC $\uparrow$ ROCAUC $\uparrow$ ROCAUC $\uparrow$ ROCAUC $\uparrow$
Sage 4.792 0.7917 0.8768 0.8424 0.7358 0.7273 0.6239 0.6590 0.6620
Gat 5.208 0.8053 0.8954 0.8428 0.7410 0.6741 0.6146 0.6638 0.6417
Pna 5.333 0.8000 0.8924 0.8464 0.7645 0.7011 0.6249 0.6378 0.6267
Hgt 5.333 0.7965 0.8805 0.8252 0.7551 0.6733 0.6260 0.5556 0.6214
DFS+MLP 6.833 0.7554 0.8248 0.7643 0.6815 0.6944 0.5456 0.5839 0.6521
DFS+Deepfm 7.625 0.7016 0.8092 0.8007 0.6667 0.7341 0.5289 0.5912 0.5844
DFS+Fttransformer 5.917 0.7473 0.8162 0.7863 0.6765 0.7412 0.5360 0.6247 0.6576
DFS+XGB 7.083 0.7600 0.8453 0.7561 0.6922 0.7219 0.5421 0.6028 0.6568
Griffin-unpretrained 3.708 0.7998 0.8941 0.8615 0.7307 0.7157 0.6246 0.6639 0.6261
Griffin-pretrained 3.042 0.8133 0.9058 0.8681 0.7417 0.7181 0.6253 0.6330 0.6468
@@
Metric ROCAUC $\uparrow$ ROCAUC $\uparrow$ ROCAUC $\uparrow$ ROCAUC $\uparrow$ ROCAUC $\uparrow$ ROCAUC $\uparrow$ ROCAUC $\uparrow$ ROCAUC $\uparrow$ RMSE $\downarrow$
Sage 0.7262 0.7554 0.6988 0.6860 0.8470 0.8558 0.8861 0.6610 0.9639
Gat 0.7299 0.7945 0.6788 0.6526 0.8284 0.8645 0.8853 0.6498 0.9563
Pna 0.7258 0.7213 0.5898 0.6558 0.8366 0.8664 0.8896 0.6684 0.9615
Hgt 0.7307 0.7226 0.5538 0.6619 0.8495 0.8670 0.8817 0.7119 0.9636
DFS+MLP 0.7134 0.7796 0.6802 0.6518 0.8181 0.8326 0.8783 0.6772 0.9847
DFS+Deepfm 0.6705 0.8133 0.6801 0.6210 0.8182 0.8212 0.8821 0.6315 0.9946
DFS+Fttransformer 0.7302 0.8250 0.6800 0.6578 0.8034 0.8376 0.8749 0.6660 0.9888
DFS+XGB 0.7158 0.8022 0.6786 0.6521 0.7906 0.8251 0.8675 0.7142 0.9972
Griffin-unpretrained 0.7052 0.7855 0.6847 0.6722 0.9512 0.8457 0.8956 0.6680 0.6117
Griffin-pretrained 0.7091 0.7795 0.6804 0.6908 0.9643 0.8435 0.8962 0.6978 0.5994
@@
Metric MAE $\downarrow$ MAE $\downarrow$ MAE $\downarrow$ MAE $\downarrow$ MAE $\downarrow$ Logloss $\downarrow$ Logloss $\downarrow$
Sage 1.3565 0.6276 1.1466 0.9191 1.5251 0.6151 2.3820
Gat 1.3695 0.6417 1.1086 0.8931 1.4098 0.6191 2.3880
Pna 0.8944 0.6084 1.4788 0.9572 1.4736 0.7015 2.3840
Hgt 0.6686 0.6212 1.5845 0.9180 1.5000 0.6126 2.3830
DFS+MLP 0.6610 0.6146 1.2518 0.9462 2.1104 0.6185 2.3840
DFS+Deepfm 1.1784 2.5796 1.3540 0.9753 2.0998 0.6118 2.3810
DFS+Fttransformer 0.6339 0.5934 0.9256 0.9793 1.6947 0.6460 2.3830
DFS+XGB 0.6743 0.6130 1.2957 0.9514 2.1859 0.6663 2.3890
Griffin-unpretrained 0.6593 0.5586 0.8879 0.7926 1.1700 0.5684 2.3930
Griffin-pretrained 0.6586 0.5694 0.8962 0.7945 1.2148 0.5518 2.3890
@@
Dataset/Task Diginetica/ctr Rel-hm/item-sales Rel-hm/user-churn Retailrocket/cvr Seznam/charge Seznam/prepay
@@
Dataset/Task Amazon/churn Amazon/rating Outbrain/ctr Rel-avito/ad-ctr Rel-avito/user-clicks Rel-avito/user-visits
@@
Dataset/Task Rel-f1/DNF Rel-f1/position Rel-f1/top3 Stackexchange/churn Stackexchange/upvote Virus/wnv
@@
Dataset/Task Airbnb/destination Rel-trial/site-success Rel-trial/study-adverse Rel-trial/study-outcome Talk/demo-pred Tel/severity
@@
Dataset/Task Diginetica/ctr Rel-hm/item-sales Rel-hm/user-churn Retailrocket/cvr Seznam/charge Seznam/prepay
@@
Dataset/Task Amazon/churn Amazon/rating Outbrain/ctr Rel-avito/ad-ctr Rel-avito/user-clicks Rel-avito/user-visits
@@
Dataset/Task Rel-f1/DNF Rel-f1/position Rel-f1/top3 Stackexchange/churn Stackexchange/upvote Virus/wnv
@@
Dataset/Task Airbnb/destination Rel-trial/site-success Rel-trial/study-adverse Rel-trial/study-outcome Talk/demo-pred Tel/severity
@@
Dataset/Task Diginetica/ctr Rel-hm/item-sales Rel-hm/user-churn Retailrocket/cvr Seznam/charge Seznam/prepay
@@
Dataset/Task Amazon/churn Amazon/rating Outbrain/ctr Rel-avito/ad-ctr Rel-avito/user-clicks Rel-avito/user-visits
@@
Dataset/Task Rel-f1/DNF Rel-f1/position Rel-f1/top3 Stackexchange/churn Stackexchange/upvote Virus/wnv
@@
Dataset/Task Airbnb/destination Rel-trial/site-success Rel-trial/study-adverse Rel-trial/study-outcome Talk/demo-pred Tel/severity
@@
Dataset/Task Diginetica/ctr Rel-hm/item-sales Rel-hm/user-churn Retailrocket/cvr Seznam/charge Seznam/prepay
@@
Dataset/Task Amazon/churn Amazon/rating Outbrain/ctr Rel-avito/ad-ctr Rel-avito/user-clicks Rel-avito/user-visits
@@
Dataset/Task Rel-f1/DNF Rel-f1/position Rel-f1/top3 Stackexchange/churn Stackexchange/upvote Virus/wnv
@@
Dataset/Task Airbnb/destination Rel-trial/site-success Rel-trial/study-adverse Rel-trial/study-outcome Talk/demo-pred Tel/severity